# DeepExtractor: Time-domain reconstruction of signals and glitches in gravitational wave data with deep learning

### v1

@@ III.2.2 Real LIGO O3 backgrounds
(a) PSD of detector noise in LIGO Hanford calculated over 512 seconds and centered at GPS time $1256665874$.
(b) PSD of the Hanford noise after whitening (blue) and simulated white noise (green) described in Section III.2.1.
Figure 6: Real vs simulated detector noise.
GW detector noise is often modeled as Gaussian and stationary over short timescales, but it is generally non-Gaussian and exhibits a time-varying PSD. The sensitivity of GW detectors is constrained by broadband noise sources, such as seismic noise at low frequencies and quantum sensing noise at high frequencies [12]. An example PSD from LIGO Hanford is shown in Figure 6. The bottom plot compares the PSD of whitened data using Welch averaging [101] to that of simulated white noise described in Section III.2. While whitening reduces broadband noise, narrow-band spectral lines—caused by electrical and mechanical sources or resonances—persist [102]. In contrast, the simulated noise PSD shows a smooth, flat trend with no spectral lines, underscoring the complexities of real GW detector noise.
@@
We restrict the scope of this study to LIGO data during O3. We accumulate real backgrounds by identifying GPS times for all GW events and Omicron triggers. Background segments of 141414\,14s are extracted from GWOSC [104] data during O3a and O3b, ensuring these segments are at least 222\,2s away from any trigger. These segments are whitened using Welch’s method [101], with the first and last 222\,2s discarded to mitigate boundary artifacts. The resulting 101010\,10s segments are divided into 222\,2s samples. We perform a final validation step using Q-scans for each $2\,\mathrm{s}$ sample, discarding any sample with a maximum pixel (tile) energy — defined as the squared magnitude of the Q-transform coefficient for a time-frequency tile — exceeds 20 (see [105] for details on pixel energy). This ensures that no sample contains excessive power that could interfere with the training of DeepExtractor. After applying this criterion, we collect a total of $45,400$ background samples from LIGO’s O3 run, distributed as follows: $18,400$ from Hanford O3a, $1,300$ from Livingston O3a, $18,600$ from Hanford O3b, and $7,100$ from Livingston O3b. [^2] While Livingston has nearly twice as many Omicron triggers as Hanford, a significant portion of Livingston background samples were also excluded during the Q-scan validation due to persistent excess power, even outside the Omicron triggers. For injections, we follow the approach described in Section III.2.1, with the key distinction that the injections are now added to real, whitened O3 backgrounds.
@@ III.2.3 Unseen Test Classes
To evaluate the generalization performance of DeepExtractor, we utilize two glitch generators: gengli and cDVGAN [106, 107]. These generative deep learning models produce realistic glitches, also enabling robust testing on unseen glitch classes. Trained on blip glitches extracted from LIGO O2 data using BayesWave, gengli generates reliable representations of this specific glitch type. A conditional GAN trained on preprocessed LIGO O3 data, cDVGAN produces blip [^3] https://gswiki.ischool.syr.edu/en/Glitches/Blip and tomte [^4] https://gswiki.ischool.syr.edu/en/Glitches/Tomte glitches along with BBH signals with component masses in the range of $[30,160]M_{\odot}$, a spin of zero and fixing $m_{1}>m_{2}$. As a conditional model, cDVGAN can create hybrid samples that blend features across all three training classes, introducing even greater diversity. By leveraging these generative models, we create a diverse test dataset that mimics the variety and complexity of real detector glitches. This dataset provides a robust benchmark for assessing DeepExtractor’s ability to generalize to unseen classes.
@@ III.3.1 Simulated Experiments
Simulated glitches are generated using both analytical waveforms and samples from the glitch generators discussed in Section III.2.3. While the analytical waveforms represent known classes from training, we include unseen classes using glitches simulated with gengli [106] and cDVGAN [107]. Samples are injected into simulated white noise at SNRs $\in[7.5,100]$ using Eq. 5. We limit the maximum SNR of the test glitches to 100 (reduced from 250 in our training set) to create a controlled environment for glitch reconstruction. Furthermore, over 90% of glitches observed during O3 had an SNR below 100.
@@ III.3.2 BayesWave Comparison
In this experiment, we inject 30 samples from each of the 12 glitch classes mentioned in Table 1 (360 samples in total) into background noise at SNRs $\in[15,100]$, then process the samples using both DeepExtractor and BayesWave for a direct, 1:1 comparison. We increase the lower SNR bound from $7.5$ (as used in Section III.3.1) to $15$ to prevent potential convergence issues with BayesWave. The performance of each method is assessed by evaluating the mismatch (Eq. 7) between the injected and reconstructed samples, following the same approach as in Section III.3.1.
III.3.3 Gravity Spy Glitches
The final experiment of this study evaluates the performance of our model on Gravity Spy glitches. We conduct a qualitative assessment of Gravity Spy Q-scans before and after glitch subtraction to visually inspect the effectiveness of our model, while also inspecting the resulting glitch and residual noise. Since our fine-tuned DeepExtractor model was trained on 222\,2s samples extracted from 141414\,14s whitened segments, we need to use 141414\,14s of real data surrounding the glitch during inference to ensure the whitening filter remains consistent (see Section III.2.2). To avoid inadvertently suppressing the glitch, we use the PSD from an adjacent 141414\,14s segment that does not include the glitch. This ensures that the glitch is not included in the PSD calculation, while still allowing us to whiten the data effectively. Although the adjacent 141414\,14s segment is not vetoed for other glitches (which could result in imperfect whitening), this approach helps prevent the desired glitch from being suppressed during the whitening process.
For a limited quantitative assessment of the model’s performance on real glitch data, we use the Shapiro-Wilk test [110]. This statistical test compares the p-values of real glitch events before and after mitigation with DeepExtractor. Samples with p-values below a confidence interval (5%) indicate strong evidence for a non-Gaussian component, which is indicative of a glitch. We present the Shapiro-Wilk results in Appendix VII.4, as this test is not a perfect method for characterizing glitches (see Section IV.3 for more details).
IV Results
IV.1 Simulated Experiments
@@
gaussian_pulse $0.4^{+0.4}_{-0.2}$ $3.8^{+3.8}_{-1.7}$ $0.6^{+0.7}_{-0.3}$ $37.7^{+13.9}_{-11.5}$ $1.8^{+2.1}_{-0.8}$ $0.6^{+0.7}_{-0.2}$ $\mathbf{0.4}^{+0.4}_{-0.2}$
ringdown $0.6^{+0.8}_{-0.2}$ $5.2^{+4.3}_{-2.1}$ $0.9^{+1.0}_{-0.4}$ $39.3^{+15.0}_{-9.5}$ $2.2^{+2.4}_{-1.0}$ $0.9^{+0.8}_{-0.4}$ $\mathbf{0.7}^{+0.6}_{-0.3}$
@@
Figure 7: DeepExtractor reconstructions (orange) of samples comprising multiple injections (black). The upper panels display the full $h(t)$ signals (gray), while the lower panels zoom in on a 0.50.50.5\,0.5s segment (outlined with a red box) to highlight interesting features of the reconstructed injections after background removal.
@@
In contrast to the Autoencoder1D model, which produces mismatch results similar to UNET1D, the Autoencoder2D model exhibits notably large mismatch values between the reconstructed and injected glitches. This suggests that feature compression within the Autoencoder layers—particularly in the spectrogram domain—may be a critical issue. To explore this further, we removed one of the Autoencoder2D layers to assess whether reduced compression could lead to better results in the spectrogram domain, but this adjustment did not improve the results beyond those presented in Table 1. This finding highlights the importance of skip connections in U-Nets, which play a vital role in preserving key features and enhancing reconstruction accuracy.
Finally, applying DeepExtractor to 512 pure background samples without injections (i.e., where it is expected to reproduce the input) yields a maximum mismatch of $0.26\%$ between the target and the reconstructed background noise. This demonstrates the model’s ability to accurately replicate the background noise in the absence of signals or glitches, ensuring that no extraneous signals or glitches are reconstructed when they are not present in the data.
FIG. 7 shows reconstructions of multi-injection samples, similar to the training data, demonstrating DeepExtractor’s versatility. It consistently recovers diverse glitches, even when heavily obscured by noise. Across 512 samples, DeepExtractor achieves a median mismatch of $0.7^{+0.2}_{-0.2}$. The model excels with multi-injection samples by effectively matching non-Gaussian components, which are more frequent per sample and better aligned with the training data than the single-class samples in Table 1555BayesWave was not applied to these samples, as it requires a single trigger time, while these samples have multiple trigger times. We instead maintain a controlled comparison with single-class samples, where results are easier to interpret..
@@ IV.2 BayesWave Comparison
FIG. 8 presents a comparison of mismatch performance between DeepExtractor and BayesWave on a reduced dataset, consisting of 30 samples per class. For reference, examples of reconstructions from both BayesWave and DeepExtractor are shown in FIG. 9. DeepExtractor consistently outperforms BayesWave, even on previously unseen test classes. Specifically, DeepExtractor achieves a median mismatch of $0.3^{+0.2}_{-0.1}\%$ across the entire test dataset, showing a significant improvement over BayesWave, which has a median mismatch of $1.9^{+2.2}_{-1.1}\%$ (the bounds represent the $1\sigma$ confidence interval). The performance improvement of DeepExtractor compared to Section IV.1 can be attributed to the lower SNR limit being raised to 15, making the reconstruction task easier.
@@
The discrepancies may be due to BayesWave’s difficulty in converging due to high-frequency features in the these samples. Furthermore, DeepExtractor demonstrates significant improvements when applied to realistic glitch samples from glitch generators, particularly hybrid samples from cDVGAN, which feature diverse glitch morphologies.
Finally, DeepExtractor also offers a substantial advantage in terms of speed, reconstructing a single glitch sample in an average of $0.14$ seconds on a CPU (or $0.05$ per glitch seconds when processing a batch of 256 samples). This represents a speedup of over $10,000$ times compared to BayesWave, which relies on a traditional Bayesian framework.
@@
Figure 9: Examples from our comparison between BayesWave and DeepExtractor. Injected SNRs and mismatches yielded from both approaches are shown above each plot.
@@ IV.3 Gravity Spy Glitches
This experiment applies transfer learning to fine-tune DeepExtractor on real LIGO O3 data from both the Hanford and Livingston detectors. The primary objective is to improve the model’s ability to reconstruct real glitch events. Given the complex, unmodeled nature of glitches, we take a qualitative approach by visualizing DeepExtractor’s reconstruction performance through Q-scan images and time-series plots.
@@
FIG. 10(d) displays the time-domain output of both the pre-trained (blue) and fine-tuned (orange) models. Without transfer learning, the model erroneously reconstructs the line features, obscuring the true glitch signal. However, after transfer learning, the model has learned to properly identify and reconstruct these features, resulting in a more accurate glitch reconstruction.
Additional examples of DeepExtractor’s fine-tuned performance on Gravity Spy glitches are provided in Appendix VII.5. The Q-scans show that, for most glitch classes, noise features are effectively preserved after glitch removal. However, for louder glitches, such as those from the ‘Extremely Loud’ and ‘Koi Fish’ classes, where the glitch dominates the signal, the removal process leads to a“zeroing out” effect on the data, as seen in the Q-scans post-mitigation. However, in most cases, the time-series plots demonstrate that the reconstructions align closely with the actual data. Overall, DeepExtractor exhibits flexibility across glitch classes.
Shapiro-Wilk test results, presented in Appendix VII.4, show that most glitch classes have average p-values above the 5% significance threshold, suggesting limitations in the Shapiro-Wilk test for characterizing glitches. However, higher p-values after glitch mitigation indicate that the residual data better approximates a Gaussian distribution, signaling that DeepExtractor has improved the data quality.

### v3

@@ III.2.2 Real LIGO O3 backgrounds
Figure 5: Power spectral density (PSD) of (a) real LIGO Hanford detector noise and (b) the same PSD whitened and our simulated white noise. In comparison, it is visible that our simulated noise misses instrumental or environmental lines commonly found in the GW detectors.
GW detector noise is often modeled as Gaussian and stationary over short timescales, but it is generally non-Gaussian and exhibits a time-varying PSD. The sensitivity of GW detectors is constrained by broadband noise sources, such as seismic noise at low frequencies and quantum sensing noise at high frequencies [17]. An example PSD from LIGO Hanford is shown in FIG. 5. The bottom plot compares the PSD of whitened data using Welch averaging [86] to that of simulated white noise described in the previous section. While whitening reduces broadband noise, narrow-band spectral lines—caused by electrical and mechanical sources or resonances—persist [87]. In contrast, the simulated noise PSD shows a smooth, flat trend with no spectral lines, underscoring the complexities of real GW detector noise.
@@
We restrict the scope of this study to LIGO data during O3. We accumulate real backgrounds by identifying GPS times for all GW events and Omicron triggers. Background segments of 141414\,14s are extracted from GWOSC [89] data during O3a and O3b, ensuring these segments are at least 222\,2s away from any trigger. These segments are whitened using Welch’s method [86], with the first and last 222\,2s discarded to mitigate boundary artifacts. The resulting 101010\,10s segments are divided into 222\,2s samples. We perform a final validation step using Q-scans for each $2\,\mathrm{s}$ sample, discarding any sample with a maximum pixel (tile) energy — defined as the squared magnitude of the Q-transform coefficient for a time-frequency tile — exceeding 20 (see [90] for details on pixel energy). This ensures that no sample contains excessive power that could interfere with the training of DeepExtractor. After applying this criterion, we collect a total of $45,400$ background samples from LIGO’s O3 run, distributed as follows: $18,400$ from Hanford O3a, $1,300$ from Livingston O3a, $18,600$ from Hanford O3b, and $7,100$ from Livingston O3b [^2] While Livingston has nearly twice as many Omicron triggers as Hanford, a significant portion of Livingston background samples were also excluded during the Q-scan validation due to persistent excess power, even outside the Omicron triggers.. For injections, we follow the approach described in Section III.2.1, with the key distinction that the injections are now added to real, whitened O3 backgrounds.
@@ III.2.3 Unseen Test Classes
To evaluate the generalization performance of DeepExtractor, we utilize two glitch generators: gengli and cDVGAN [91, 92]. These generative deep learning models produce realistic glitches, also enabling robust testing on unseen glitch classes. Trained on blip [^3] https://gswiki.ischool.syr.edu/en/Glitches/Blip glitches extracted from LIGO O2 data using BayesWave, gengli generates reliable representations of this specific glitch type in both Hanford and Livingston detectors separately. A conditional GAN trained on preprocessed LIGO O3 data, cDVGAN produces blip and tomte [^4] https://gswiki.ischool.syr.edu/en/Glitches/Tomte glitches along with BBH signals with component masses in the range of $[30,160]M_{\odot}$, a spin of zero and fixing $m_{1}>m_{2}$. As a conditional model, cDVGAN can create hybrid samples (simplex and uniform) that blend features across all three training classes, introducing even greater diversity. By leveraging the class-mixing feature of cDVGAN, we create a diverse test dataset to assess DeepExtractor’s ability to generalize to unseen classes. Examples of gengli and cDVGAN generations can be found in Appendix VII.3.
@@ III.3.1 Simulated Experiments
Simulated glitches are generated using both analytical waveforms and samples from the glitch generators discussed in the previous section. While the analytical waveforms represent known classes from training, we include unseen classes using glitches simulated with gengli and cDVGAN. Samples are injected into simulated white noise at SNRs $\in[7.5,100]$ using Eq. 3. We limit the maximum SNR of the test glitches to 100 (reduced from 250 in our training set) to create a controlled environment for glitch reconstruction. Furthermore, over 90% of high-confidence glitches [^5] Those classified with a Gravity Spy confidence of at least 90%. observed during O3 had an SNR below 100.
@@ III.3.2 BayesWave Comparison
In this experiment, we inject 30 samples from each of the 12 glitch classes mentioned in Table 1 (360 samples in total) into background noise at SNRs $\in[15,100]$, then process the samples using both DeepExtractor and BayesWave for a direct, 1:1 comparison. We increase the lower SNR bound from $7.5$ (as used in Section III.3.1) to $15$ to prevent potential convergence issues with BayesWave. The performance of each method is assessed by evaluating the mismatch (Eq. 5) between the injected and reconstructed samples, following the same approach outlined in the previous section.
@@
gaussian_pulse $\mathbf{0.4}^{+0.4}_{-0.2}$ $3.8^{+3.8}_{-1.7}$ $0.6^{+0.7}_{-0.3}$ $37.7^{+13.9}_{-11.5}$ $1.8^{+2.1}_{-0.8}$ $0.6^{+0.7}_{-0.2}$ $\mathbf{0.4}^{+0.4}_{-0.2}$
ringdown $\mathbf{0.6}^{+0.8}_{-0.2}$ $5.2^{+4.3}_{-2.1}$ $0.9^{+1.0}_{-0.4}$ $39.3^{+15.0}_{-9.5}$ $2.2^{+2.4}_{-1.0}$ $0.9^{+0.8}_{-0.4}$ $0.7^{+0.6}_{-0.3}$
@@
III.3.3 Gravity Spy Glitches
This experiment evaluates the performance of our model on Gravity Spy glitches. We conduct a qualitative assessment of Gravity Spy Q-scans before and after glitch subtraction to visually inspect the effectiveness of our model, while also inspecting the resulting glitch reconstruction in the time domain. Since our fine-tuned DeepExtractor was trained on 222\,2s samples extracted from 141414\,14s whitened segments, we need to use 141414\,14s of real data surrounding the glitch during inference to ensure the whitening filter remains consistent (see Section III.2.2). To avoid inadvertently suppressing the glitch during whitening, we estimate the PSD used for whitening from an adjacent 141414\,14s segment via Welch’s method that does not include the glitch. This approach ensures the glitch does not contaminate the PSD estimate, while still enabling effective whitening of the data. Although the adjacent 141414\,14s segment is not vetoed for other glitches (which could result in imperfect whitening), this approach helps prevent the desired glitch from being suppressed during the whitening process.
III.3.4 Reconstructing O3 signals
To showcase the potential of DeepExtractor for reconstructing real gravitational wave signals in a model-agnostic way, we apply it to three events from the O3 observing run: GW190521_074359, GW200129_065458, and GW200224_222234. The component masses for these events were, respectively, 45.4 $M_{\odot}$ and 33.4 $M_{\odot}$, 34.5 $M_{\odot}$ and 29.0 $M_{\odot}$, and 40.0 $M_{\odot}$ and 32.7 $M_{\odot}$. These are among the loudest events reported during O3 that were confidently detected by both the Hanford and Livingston detectors. We follow the same preprocessing pipeline as in the previous section: whitening a 141414\,14s segment centered around the event, and extracting a 222\,2s window around the merger, with the merger time aligned at 1.851.851.85\,1.85s within the segment. Similarly, we calculate the PSD from the previous 141414\,14s segment to whiten the data, to ensure that the signal itself is not suppressed during the whitening process. We then compare DeepExtractor’s reconstructions to the maximum likelihood templates obtained from parameter estimation, once again using the time-domain mismatch as the evaluation metric. The template itself is whitened using the same PSD as above, since the original template is given according to coloured detector noise.
IV Results
IV.1 Simulated Experiments
Figure 6: DeepExtractor reconstructions (orange) of samples comprising multiple injections (black). The upper panels display the full 222\,2s of whitened strain (gray), the middle panels provide a zoomed-in view of the region outlined in red, emphasizing key features of the reconstructions and the bottom panels display the residual between the reconstructed and injected glitches.
@@ IV.1 Simulated Experiments
In contrast to the Autoencoder1D model, which produces mismatch results similar to UNET1D, the Autoencoder2D model exhibits notably large mismatch values between the reconstructed and injected glitches. This suggests that feature compression within the Autoencoder layers—particularly in the spectrogram domain—may be a critical issue. To explore this further, we removed one of the Autoencoder2D layers to assess whether reduced compression could lead to better results in the spectrogram domain, but this adjustment did not improve the performance. This finding highlights the importance of skip connections in U-Nets, which play a vital role in preserving key features and enhancing reconstruction accuracy.
Finally, applying DeepExtractor to 512 pure background samples without injections (i.e. where it is expected to reproduce the input) yields a maximum mismatch of $0.26\%$ between the target and the reconstructed background noise. This demonstrates the model’s ability to accurately replicate the background noise in the absence of signals or glitches, ensuring that no extraneous signals or glitches are reconstructed when they are not present in the data.
FIG. 6 shows reconstructions of multi-injection samples, similar to the training data, demonstrating DeepExtractor’s versatility. Across 512 of such samples, DeepExtractor achieves a median mismatch of $0.7^{+0.2}_{-0.2}$. The model excels with multi-injection samples by effectively matching non-Gaussian components, which are more frequent per sample and better aligned with the training data than the single-class samples in Table 1666BayesWave was not applied to these samples, as it requires a single trigger time, while these samples have multiple trigger times. We instead maintain a controlled comparison with single-class samples, where results are easier to interpret..
@@ IV.2 BayesWave Comparison
FIG. 7 presents a comparison of mismatch performance between DeepExtractor and BayesWave on a reduced dataset, consisting of 30 samples per class. Examples of reconstructions from both BayesWave and DeepExtractor are shown in FIG. 8. DeepExtractor consistently outperforms BayesWave, even on previously unseen test classes. Specifically, DeepExtractor achieves a median mismatch of $0.3^{+0.2}_{-0.1}\%$ across the entire test dataset, showing a significant improvement over BayesWave, which has a median mismatch of $1.9^{+2.2}_{-1.1}\%$ (the bounds represent the $1\sigma$ confidence interval). The performance improvement of DeepExtractor compared to the previous section can be attributed to the lower SNR limit being raised to 15, making the reconstruction task easier.
@@
The discrepancies may be due to BayesWave’s difficulty in converging due to high-frequency features in these samples. Furthermore, DeepExtractor demonstrates significant improvements when applied to realistic glitch samples from glitch generators, particularly hybrid samples from cDVGAN, which feature diverse glitch morphologies.
Finally, DeepExtractor also offers a substantial advantage in terms of speed, reconstructing a single glitch sample in an average of $0.14$ seconds on a CPU (or $0.05$ seconds per glitch when processing a batch of 256 samples). This represents a speedup of over $10,000$ times compared to BayesWave.
The significantly longer runtime of BayesWave stems from its use of RJMCMC to explore a multidimensional parameter space and to construct posterior distributionss. While computationally intensive, this Bayesian approach provides uncertainty estimates and flexible modeling of both glitches and astrophysical signals with minimal assumptions.
In contrast, DeepExtractor performs fast, deterministic inference via a single forward pass through a trained neural network, bypassing the need for iterative sampling or likelihood evaluation. This makes it well-suited for real-time or large-scale applications. While it trades some flexibility and uncertainty quantification for speed, DeepExtractor offers rapid, high-fidelity reconstructions of glitches in scenarios where fast turnaround is critical or full Bayesian inference is computationally prohibitive.
@@
Figure 8: Examples from our comparison between BayesWave and DeepExtractor. Injected SNRs and mismatches yielded from both approaches are shown above each plot. The bottom panels display the residual between the reconstructed and injected glitches for BayesWave and DeepExtractor.
@@ IV.3 Gravity Spy Glitches
Figure 10: Reconstructions for three examples from distinct Gravity Spy glitch classes: Light Modulation, Low Frequency Burst, and Paired Doves. The top row displays Q-scans of the input to the network. The middle row shows Q-scans of the residual after subtracting DeepExtractor’s reconstruction. The bottom row presents the corresponding time-domain input and reconstructed waveforms. The maximum color limit in the Q-scans is set to 25, similarly to the Q-scans shown in FIG. 9.
This experiment applies transfer learning to fine-tune DeepExtractor on real LIGO O3 data from both the Hanford and Livingston detectors. The primary objective is to improve the model’s ability to reconstruct real glitch events. Given the complex, unmodeled nature of glitches, we take a qualitative approach by visualizing DeepExtractor’s reconstruction performance through Q-scans and time-series plots.
@@
FIG. 9d displays the time-domain output of both the pre-trained and fine-tuned models. Without transfer learning, the model erroneously reconstructs the line features, obscuring the true glitch signal. However, after transfer learning, the model compensates for these features, resulting in a more accurate glitch reconstruction.
FIG 10 presents examples of DeepExtractor reconstructions following transfer learning, illustrating performance across three distinct glitch classes. Additional reconstruction examples are provided in Appendix VII.4. The Q-scans show that, for most glitch classes, noise features are effectively preserved after glitch removal. However, for louder glitches, such as those from the ‘Extremely Loud’ and ‘Koi Fish’ classes, where the glitch dominates the sample, the removal process leads to a ‘zeroing out’ effect on the data, as seen in the Q-scans post-mitigation. However, in most cases, the time-series plots demonstrate that the reconstructions align closely with the actual data. Overall, DeepExtractor exhibits flexibility across glitch classes.
IV.4 Reconstructing O3 signals
In this section, we present the reconstructions of three gravitational-wave events detected during O3 observing run: GW190521_074359, GW200129_065458, and GW200224_222234. DeepExtractor is applied independently to whitened strain data from each detector around the time of merger.
FIG. 11 shows the reconstructed signals for all three events in both LIGO detectors. For each case, we report the SNR, the mismatch relative to the maximum likelihood waveform [96], and the residual between the reconstructed and the maximum likelihood waveform.
Although DeepExtractor was not explicitly trained on astrophysical signals and is agnostic to their morphologies, it achieves mismatches below 10% for most cases. The exception is the GW190521_074359 event in the Hanford detector, where the mismatch reaches 16.45%.
This degradation in performance compared to the simulated setting is likely due to a misalignment between our preprocessing pipeline and that of the parameter estimation pipeline—specifically our whitening approach for both strain data and parameter estimation templates, which is necessary for input to DeepExtractor. Improving the alignment of DeepExtractor’s training scheme with parameter estimation pipelines and incorporating signal models into training could enhance the performance.
This experiment demonstrates the promise of DeepExtractor as a model-agnostic tool for reconstructing gravitational-wave signals. In cases where glitches are coincident with signals, DeepExtractor reconstructs both the signal and glitch simultaneously—a scenario we also explore in Appendix VII.5. Future work will involve extending DeepExtractor to separate both GWs and glitches during reconstruction (see Section V for more discussion).
(a) GW190521_074359
(b) GW200129_065458
(c) GW200224_222234
Figure 11: Reconstruction of gravitational-wave signals from three O3 events using DeepExtractor. Each panel shows the comparison between the reconstructed waveform, the template, and the residuals for H1 and L1 detectors. The matched-filter SNR ($\rho_{MF}$) in each detector frame is shown at the top of the respective plots, along with the mismatch ($\mathcal{M}$) between the reconstruction and the maximum likelihood template. We acknowledge that misalignment between our data preprocessing and that of the parameter estimation pipeline likely contributes to the observed mismatches.